# AutoInst: Automatic Instance-Based Segmentation of LiDAR 3D Scans

### v1

@@ I Introduction
Operation of autonomous systems in real outdoor conditions involves continuously interpreting sensor data representing their 3D environment; commonly, an initial stage involves 3D segmentation, constructing a semantics- or instance-aware environment map, ideally defined in any spatial location and at each step in time. Making use of progress in sensing equipment, panoptic 3D LiDAR segmentation [1, 2] conveniently provides such a holistic environment description by densely assigning semantic and instance labels to individual 3D points. Recently, large collections of 3D point clouds with dense, high-quality annotations have enabled training effective panoptic 3D segmentation methods [3, 4, 5] powered by deep neural networks (NNs).
@@
Unfortunately, obtaining and, most importantly, annotating such datasets is significantly time- and labour-consuming [6, 7, 8, 9] as 3D objects are difficult to recognize and accurately segment in sparse and noisy 3D LiDAR scans. Recognizing this, recently proposed 3D instance segmentation approaches [10, 11, 12, 13] seek to minimize dependence on annotated collections by integrating advanced machine learning methods. Naively using neural models optimized on existing labeled collections [14], however, is suboptimal due to variations in sensor configurations and acquisition conditions across datasets. Despite this mismatch in data distributions being addressed by domain adaptation methods [15] to some extent, automatically adding new semantic classes remains impossible. Approaches based on active or weakly-supervised learning [16, 17] enable using sparse or incomplete labels but commonly need user input to guide initial predictions. Very recently, distilling segmentation foundational models has seen interest for indoor domain [18, 19, 14, 20] but remains in its infancy for outdoor LiDAR data [21]. In this work, we take an alternative path and develop an automatic learning-based algorithm for instance-based segmentation of 3D LiDAR point clouds, requiring no semantic or instance annotations during training and focusing on dense aggregated scans. As a result, instance segmentations produced by our method are class-agnostic; to implement various downstream tasks, they can be mapped to respective class nomenclatures and instance granularities. We make two key design decisions leading to important implications for our method. First, instead of processing single scans as done by most existing approaches [22, 1, 3], we follow the common practice for manual labeling [6] and operate with dense registered 3D LiDAR acquisitions to facilitate our algorithm to produce universally useful, high-quality instance segmentations. By improving the sampling density of the scanned 3D environment, this enables segmenting instances with smaller linear dimensions (e.g., vegetation or traffic signs) more reliably. We generally choose to produce finer granularity instances at a given sampling density, as merging these is significantly easier compared to splitting. Dense 3D point clouds additionally benefit dense image-to-point correspondences, enabling effective use of image-based features, but require addressing several important issues. We exclude dynamic objects (such as moving cars or humans) which lead to noisy instances, focusing on segmenting static environments instead. To enable processing spatially extended 3D environments represented by billions of data points [6, 23], we predict instances in local overlapping point chunks and design a scalable merging algorithm for fusing predictions obtained in each chunk. As our second design choice, drawing inspiration from the recent line of work targeting image and 3D indoor domains [24, 25, 26], we design our framework to operate in two stages: unsupervised generation and self-trained refinement of mask proposals. For generation of our instance mask proposals, we integrate recently proposed deep self-supervised point-[27] and image-based [28] representations constructed from aligned 3D LiDAR and image captures, providing dense 3D features with high predictive power. Combining these descriptors with spatial prior results in a weighted proxy-graph connecting scanned 3D points; cutting this graph [29] produces initial, possibly noisy, instance segmentations. To eliminate spurious or noisy instances and maximize segmentation performance, we refine initial mask proposals by performing self-training using a procedure similar to [24, 25, 26]. We evaluate our algorithm on SemanticKITTI [6], mapping our segmentations to ground-truth instances. To quantify advantages of our learning-based algorithm, we compare it to non-learnable [30, 19], and learnable unsupervised [22] and supervised [3] instance segmentation baselines. We assess contributions of individual components and identify an optimal configuration of features for learning-based instance generation and refinement. We contribute:
@@ II Related Work
Panoptic [1, 2] and Instance [31] 3D Segmentation
these are key 3D computer vision tasks with vast applications in autonomous robotics, where they serve to partition an input point cloud into distinct object instances and (panoptic only) background regions. Earlier approaches produce instances by spatially clustering [32, 30] segmentations predicted by pretrained scene understanding models but crucially depend on the performance of the backbone [33, 34]. For realtime operation of mobile robots, learning-based methods have been proposed for panoptic [1, 35, 36, 37, 38, 4, 3, 39, 40, 5] and instance-based [31, 41] segmentation of single-scan 3D acquisitions. Alternatively, segmentation of dense 3D point clouds (similar to this work) can aid annotation of 3D data or perform scene analytics. Here, static indoor scenes have seen the most progress with the emergence of panoptic [42, 43, 44] and instance [45, 46, 47, 48, 49] understanding schemes; however instance-based segmentation of dense outdoor LiDAR 3D data remains almost unexplored [50]. All these approaches are supervised and require large-scale annotated 3D LiDAR datasets [6, 23]. We adopt a recent architectural solution [3] in our proposal refinement network.
@@
Multiple works construct and utilize sparse annotations (e.g., a single click or a scribble per instance) [17, 16], or perform self-supervised pre-training of representations coupled with supervised fine-tuning for respective tasks [27, 51]. Still, they predominantly target either only semantic segmentation [51, 17, 16] or only indoor scenes [52, 53, 54, 55]. A recent domain adaptation method [15] transfers a pre-trained panoptic model between source and target domains, minimizing gap between them. None of these methods can handle new object types emerging during the inference stage. Weakly- and semi-supervised methods can to some extent reduce the requirements for annotations but are unable to exclude them entirely. We integrate self-supervised point embeddings [27] in our instance proposal stage.
@@
To avoid the need for annotated 3D data and enable integrating object types unseen during training, multiple approaches opt to generate and refine 3D instance proposals without associated semantic labels (i.e., class-agnostic). Many such approaches originate from clustering algorithms and use various point cloud representations including raw 3D points [56], radial projections and range images [57, 58, 59, 60], voxel grids [61, 62], curved voxels [63], etc. Recently, unsupervised pointwise features [51] were explored as a starting representation for clustering [22]. To ease segmentation of above-ground objects, several methods start with ground extraction [57, 60, 64]; likewise, our approach integrates a robust ground estimator [65] for this purpose. We compare to a recent unsupervised method [22] in our experiments. For the indoor domain, recent methods [25, 26] propose generating 3D instance proposals using spatial proximity and color similarity, and refining proposed masks by a self-trained neural network [48]; our work is conceptually similar but targets outdoor 3D LiDAR data. Building on the advanced representation capabilities of CLIP [66], open-vocabulary semantic [10] and instance [11] segmentation of 3D indoor scans were proposed. Multiple recent approaches [19, 18, 14] distill 2D segmentation foundational models [67], transferring instances predicted in a set of images to 3D space. Among these, we compare to a simple mask merging baseline [19].
@@ III-A Scan Preprocessing and Chunk Extraction
We assume that a set of LiDAR 3D scans $\{P_{t}\}_{t=1}^{T}$ with their world 3D poses $\{T_{t}^{\text{P}}\}_{t=1}^{T}$, as well as a set of images $\{I_{t}\}_{t=1}^{T}$ and their world 3D poses $\{T_{t}^{\text{I}}\}_{t=1}^{T}$ are produced by a scanning system including one or more LiDAR sensors and RGB cameras. Unless provided, accurate 3D poses can be obtained using GNSS reference poses or SLAM [68]. Operating with individual 3D LiDAR scans [22] complicates scene understanding tasks as such data rarely has high sampling density [69]; instead, to enable reliable 3D instance segmentation, we compute a dense 3D point cloud. For this, we transform 3D LiDAR and image data to a shared reference frame; dynamic objects can be detected and eliminated by recent effective methods [70]. During this aggregation step, we perform ground segmentation [65], effectively removing points that do not contain instances and allowing to exploit the geometric separation between objects. To reduce resource constraints, we additionally perform voxel downsampling with a voxel size of 5 cm. To enable using image-based features, we build an inverted index associating each point $\bm{p}$ to a set of pairs $A_{\bm{p}}=\{(I_{k},\bm{u}_{k})\}$ where $\bm{u}_{k}$ is a pixel in view $I_{k}$ capturing $\bm{p}$. To prevent points occluded in a given view from being assigned its pixels, we identify occluded points w.r.t. each view [71]. For our large-scale scenes, processing entire LiDAR 3D point maps is computationally infeasible as they include billions of sampled points; instead, we extract local overlapping point chunks for all subsequent steps. Our chunks are $w\times h\times d$ axis-aligned rectangular parallelepipeds, sampled along the cars trajectory to maximize map coverage. In our experiments, we evaluate the influence of chunk size on performance.
@@ IV-A Experimental Setup
[…]
@@
We set $\theta^{\text{I}}=0.1$, $\theta^{\text{P}}=0.5$ for S+P+I. For S+P, the point feature weight is set to $\theta^{\text{P}}=0.5$, all of our configurations use a spatial weight of $\theta^{\text{S}}=1$. We use the cutting thresholds as mentioned in Sec. III-B. To generate the final instance masks, we run inference on downsampled input dense chunks and transfer the predicted labels to the denser point cloud by nearest-neighbor assignment. We study the impact of parameter choices in Sec. IV-C.
@@
We define two unsupervised and two supervised baselines. HDBSCAN [30] is a density-based clustering algorithm, while 3DUIS [22] is an unsupervised instance segmentation method, refining HDBSCAN instance proposals with deep point-based features [51] and graph cuts. SAM3D [19] is an algorithmic method merging 2D SAM [18] predictions into 3D instance masks. We pre-process 2D SAM masks by rejecting smaller sub-instances using an IoU threshold of 0.9 and perform hidden point removal. Additionally, we train supervised MaskPLS [3], a state-of-the-art 3D panoptic segmentation method. For experiments in Tab. II, we train on sequences 0–7 and 9, holding out 8 and 10 as the test set. Further experiments (see Tab. I and Tab. III), do not involve supervised training, and were tested on all the sequences.
@@ IV-B Comparisons to State-of-the-Art
We start by evaluating our method against baselines; we present quantitative evaluation results against the unsupervised baselines in Tab. I and a quantitative comparison against the supervised MaskPLS method [3] in Tab. II. Our method (NCut(S + P) + Refine), the best variant according to ablations IV-C, outperforms unsupervised baselines (HDBSCAN and 3DUIS), as well as SAM3D that uses a supervised network. While our approach shows a moderate decrease w.r.t. the fully supervised MaskPLS, it does not require manual labeling and can be used to speed up the annotation for supervised training scenarios. Qualitative comparisons are presented in Fig. 5 for baselines and our algorithm without and with refinement step. One noticeable effect of refinement is that it splits instances that are merged together in initial proposals.
@@ IV-C Ablative Studies
To identify the optimal ingredients in our instance mask proposal step, we compare applying NCut to point graphs weighted by various combination of features, without any refinement, and present the results in Tab. III (see notation in its caption). We conclude that point embeddings (P) are crucial for obtaining strong performance; in contrast (and to our surprise), image embeddings (I) noticeably decrease performance across measures. To understand whether an alternative proposal generation step may lead to obtaining comparably strong results, we generate instance mask proposals by HDBSCAN and refine them using our self-training step. This did not lead to an improvement in accuracy of instance masks.
@@
[…]

### v2

@@ I Introduction
Operation of autonomous systems in real outdoor conditions involves continuously interpreting sensor data representing their 3D environment; commonly, an initial stage involves constructing a holistic 3D environment description by assigning semantic and instance labels to individual 3D points. Typically, neural networks (NNs) [1, 2] are trained for this task, requiring access to a large, densely annotated dataset, which is widely known to be costly and laborious to obtain [3, 4, 5]. Using neural models optimized on existing labeled collections [6], however, is suboptimal due to variations in sensor configurations and acquisition conditions across datasets, even when domain adaptation methods [7] are employed.
@@
Approaches based on active or weakly-supervised learning [8] enable using sparse or incomplete labels but commonly need user input to guide initial predictions. Very recently, distilling segmentation foundational models has seen interest for the indoor domain [9, 10, 6] but remains in its infancy for outdoor LiDAR data [11]. In this work, we take an alternative path, proposing an automatic learning-based algorithm for instance-based segmentation of dense aggregated 3D LiDAR point clouds, which requires no semantic or instance annotations during training and produces class-agnostic segmentations that can be mapped to various class nomenclatures and instance granularities for downstream tasks. We make two key design decisions leading to important implications for our method. First, instead of processing individual scans as done by most existing approaches [12, 1], we follow the manual labeling approach [3] and construct dense registered 3D LiDAR acquisitions to improve the quality of our instance segmentations. By improving the sampling density, this allows us to reliably segment smaller objects and exploit dense image-to-point correspondences, enabling effective use of image-based features. As our second design choice, drawing inspiration from the 3D indoor domain [13, 14, 15], we design our framework to operate in two stages: unsupervised generation and self-trained refinement of mask proposals. For generation of our instance mask proposals, we integrate recent deep self-supervised point-[16] and image-based [17, 18] representations constructed from aligned LiDAR and image captures, providing dense 3D features with high predictive power. We combine these 3D descriptors with spatial priors in a weighted proxy-graph and perform normalized cuts [19] obtaining initial, possibly noisy, instance segmentations. To eliminate spurious or noisy instances and maximize segmentation performance, we refine mask proposals by performing self-training using a procedure similar to [14, 15]. We evaluate our algorithm on SemanticKITTI [3], mapping our segmentations to ground-truth instances. To quantify advantages of our learning-based algorithm, we compare it to non-learnable [20, 10], and learnable unsupervised [12], as well as supervised [1] [10] instance segmentation baselines. We assess contributions of our components and identify an optimal configuration of features for learning-based instance generation and refinement. We contribute:
@@ II Related Work
Panoptic and Instance 3D Segmentation
Panoptic [21] and instance [22] 3D LiDAR segmentation are key 3D computer vision tasks with vast applications in autonomous robotics, where they serve to partition an input point cloud into distinct object instances and background regions. Earlier approaches produce instances by spatially clustering [23, 20] segmentations predicted by pretrained scene understanding models but depend on the performance of the backbone [24, 25]. For realtime operation of mobile robots, learning-based methods have been proposed for panoptic [1] and instance-based [22] segmentation of single-scan 3D acquisitions. Alternatively, segmentation of dense 3D point clouds (similar to this work) can aid annotation of 3D data or perform scene analytics. Here, static indoor scenes have seen the most progress with the emergence of panoptic [26, 27] and instance [28] understanding schemes; however instance-based segmentation of dense outdoor LiDAR 3D data remains almost unexplored [29]. All these approaches are supervised and require large-scale annotated 3D LiDAR datasets [3, 30].We adopt a recent architectural solution [1] in our proposal refinement network.
@@
Multiple works construct and utilize sparse annotations (e.g., a single click or a scribble per instance) [8], or perform self-supervised pre-training of representations coupled with supervised fine-tuning for respective tasks [16, 31]. Still, they predominantly target either only semantic segmentation [31, 8] or only indoor scenes [32]. A recent domain adaptation method [7] transfers a pre-trained panoptic model between source and target domains, minimizing the gap between them. None of these methods can handle new object types emerging during the inference stage. Weakly- and semi-supervised methods can to some extent reduce the requirements for annotations but are unable to exclude them entirely. We integrate self-supervised point embeddings [16] in our instance proposal stage.
@@
To avoid the need for annotated 3D data and enable integrating object types unseen during training, multiple approaches opt to generate and refine 3D instance proposals without associated semantic labels (i.e., class-agnostic). Many such approaches originate from clustering algorithms and use various point cloud representations including raw 3D points [20], radial projections, range images [33] and voxel grids [34]. Recently, unsupervised pointwise features [31] were explored as a starting representation for clustering [12]. To ease the segmentation of above ground objects, recent methods leverage ground segmentation [33, 12]; likewise, our approach integrates a robust ground estimator [35] for this purpose. For the indoor domain, recent methods [14, 15] generate 3D instance proposals using deep self-supervised and traditional features, refining proposed masks by a self-trained neural network [28]; our work is conceptually similar but targets outdoor 3D LiDAR data. Multiple recent approaches [10, 9, 6] distill labels from 2D segmentation foundational models [18] into 3D space. Among these, we compare to a simple mask merging baseline [10].
@@ III-A Scan Preprocessing and Chunk Extraction
We assume that a set of LiDAR 3D scans $\{P_{t}\}_{t=1}^{T}$ with their world 3D poses $\{T_{t}^{\text{P}}\}_{t=1}^{T}$, as well as a set of images $\{I_{t}\}_{t=1}^{T}$ and their world 3D poses $\{T_{t}^{\text{I}}\}_{t=1}^{T}$ are produced by a scanning system including one or more LiDAR sensors and RGB cameras. Unless provided, accurate 3D poses can be obtained using GNSS reference poses and further refined with SLAM [36], due to the synchronization of sensors, adjacent LiDAR scans and image captures can be obtained for each 3D pose in the global map. We further obtain the relevant camera features by projecting the pixels onto the aggregated pointcloud and leverage hidden-point removal. Operating with individual 3D LiDAR scans [12] complicates scene understanding tasks as such data rarely has high sampling density [37]; instead, to enable reliable 3D instance segmentation, we compute a dense 3D point cloud. For this, we transform 3D LiDAR and image data to a shared reference frame; dynamic objects can be detected and eliminated by recent effective methods [38]. During this aggregation step, we perform ground segmentation [35], effectively removing points that do not contain instances and allowing to exploit the geometric separation between objects. To reduce memory constraints, we additionally perform voxel downsampling with a voxel size of 5 cm. Due to the high density of the resulting pointcloud, our method as well the baselines operate on a further down-sampled pointcloud, with predictions being reprojected onto the dense map with 5 cm voxel size for evaluation. To enable using image-based features, we build an inverted index associating each point $\bm{p}$ to a set of pairs $A_{\bm{p}}=\{(I_{k},\bm{u}_{k})\}$ where $\bm{u}_{k}$ is a pixel in view $I_{k}$ capturing $\bm{p}$. To prevent points occluded in a given view from being assigned its pixels, we identify occluded points w.r.t. each view [39]. For our large-scale scenes, processing entire LiDAR 3D point maps is computationally infeasible as they include billions of sampled points; instead, we extract local overlapping point chunks for all subsequent steps. Our chunks are $w\times h\times d$ axis-aligned rectangular parallelepipeds, sampled along the cars trajectory to maximize map coverage. In our experiments, we evaluate the influence of chunk size on performance.
@@ IV-A Experimental Setup
We experimentally evaluate our algorithm using SemanticKITTI [3] which provides LiDAR 3D scans with instance annotations and registered RGB images acquired by a moving car in outdoor scenes, obtaining scenes on the order of $10^{9}$ points after aggregation. We use sequences 0–10 providing test labels, except sequences 1 and 4 that predominantly consist of dynamic instances. For accurate global alignment of LiDAR 3D scans, we use 3D poses computed using a surfel-based SLAM approach [36]. During mask proposal and refinement stages, we operate on 25$\times$25$\times$25 m3 point chunks; at inference time, we sample chunk centers along vehicle trajectory spaced 22 m apart (3 m overlap), resulting in 30 chunks per 1K raw LiDAR scans on average. We extract image-based features from the images obtained by the left camera. For training our refinement network and the supervised baseline, we sample point chunks, our proposals, and ground-truth labels, from KITTI sequences 0–10 with a larger 24 m overlap, obtaining a dataset with 14K chunks. To reduce memory consumption, we uniformly subsample each chunk to roughly $60$K points. We train our refinement NN [1] for 7 epochs ($12$ h) on RTX 4090 GPU using Adam [46] with a learning rate of $10^{-4}$.
@@
We set $\theta^{\text{I}}=0.1$, $\theta^{\text{P}}=0.5$ for S+P+I. For S+P, the point feature weight is set to $\theta^{\text{P}}=0.5$, all of our configurations use a spatial weight of $\theta^{\text{S}}=1$. We use the cutting thresholds as mentioned in Sec. III-B. To generate the final instance masks, we run inference on downsampled input chunks and transfer the predicted labels to the denser point cloud by nearest-neighbor assignment. We study the impact of parameter choices in Sec. IV-C.
@@
We define two unsupervised and two supervised baselines. HDBSCAN [20] is a density-based clustering algorithm, while 3DUIS [12] is an unsupervised instance segmentation method, refining HDBSCAN instance proposals with deep point-based features [31] and graph cuts. SAM3D [10] is a method merging 2D SAM [9] predictions into 3D instance masks. We pre-process 2D SAM masks by rejecting smaller sub-instances using an IoU threshold of 0.9 and perform hidden point removal. Additionally, we train supervised MaskPLS [1], a state-of-the-art 3D panoptic segmentation method. For experiments in Tab. II, we train on sequences 0–7 and 9, holding out 8 and 10 as the test set. Further experiments (see Tab. I and Tab. III), were tested on all sequences and do not include supervised training.
@@ IV-B Comparisons to State-of-the-Art
We start by evaluating our method against baselines; we present quantitative evaluation results against the unsupervised baselines in Tab. I and a quantitative comparison against the supervised MaskPLS method [1] in Tab. II. Our best-performing method (NCut(S + P) + Refine) outperforms point-based methods HDBSCAN and 3DUIS and the supervised method SAM3D without using image-based features. We refer to our ablations IV-C for a comparison of our point-based method (Ncuts(S+P)) with a multi-modal implementation (NCut(S + P + I)). While our approach shows a moderate decrease w.r.t. the fully supervised MaskPLS, it does not require manual labeling and can be used to speed up the annotation for supervised training scenarios. Qualitative comparisons are presented in Fig. 5 for baselines and our algorithm without and with refinement step. One noticeable effect of refinement is that it splits instances that are merged together in initial proposals.
@@ IV-C Ablative Studies
To identify the optimal ingredients in our instance mask proposal step, we compare applying NCut to point graphs weighted by various combination of features, without any refinement, and present the results in Tab. III (see notation in its caption). We conclude that point embeddings (P) are crucial for obtaining strong performance; in contrast (and to our surprise), image embeddings (I) noticeably decrease performance across measures. To understand whether an alternative proposal generation step may lead to obtaining comparably strong results, we generate instance mask proposals by HDBSCAN and refine them using our self-training step. While this improves upon HDBSCAN, our method achieves better results.
@@
We assess the influence of chunk and voxel size used by evaluating the performance and runtime of our method, these ablations are performed on a single sequence of KITTI [3]. Results presented in Tabs. IV and V indicate that, for our configuration and dataset, our default chunk size (25$\times$25$\times$25 m3) and voxel size (0.35 m) are close to optimal. Larger voxel size leads to coarser instances while finer one yields noisier instances. We additionally made an effort to find an optimal combination of parameters $\theta^{\text{S}},\theta^{\text{P}},\theta^{\text{I}}$ in (2). From Fig. 4 we conclude that our default values lead our proposals to match the organic instance distributions.